Condense the expression.
1/3 log5 8-1/2 log5 x = 4 log5 y

Answer by Alan3354(21583)

Condense the expression.
1/3 log5 8-1/2 log5 x = 4 log5 y
-----------------
1/3 log(8) = log(2) in any base
log(2) - log(sqrt(x)) = log(y^4) all logs to base 5
log(2/(sqrt(x)) = log(y^4)
log(2*sqrt(x)/x) = log(y^4)
2sqrt(x)/x = y^4
2sqrt(x) = xy^4